OpenStudy (anonymous):

Suppose the population of a town is3,400 in 2000. The population decreases at a rate of 2% every 20 years. What will be the projected population in 2040? Round your answer to the nearest whole number. So would the answer be 3265.36 ?

OpenStudy (whpalmer4):

population is 3400 in 2000. After 20 years (taking us to 2020), 2% of the population goes away, leaving 98% of the existing population, or 3400*0.98 = 3332. In another 20 years, it will be 2040. Another 2% will have left. What will the population be?

OpenStudy (anonymous):

I dont know?

OpenStudy (whpalmer4):

Did you see how I calculated the population decline for the first 20 years? Just apply that same process again for the next 20 years.

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

"Suppose the population of a town is3,400 in 2000. The population decreases at a rate of 2% every 20 years. What will be the projected population in 2040? Round your answer to the nearest whole number." I'd strongly suggest using an exponential growth/decay model, which looks like y=Ae^(rt). In this case, A would represent the initial population (that is, the population 3,400 in the year 2000, which would be represented by t=0; r would be the rate of decrease, which here would be written as r=-0.02; t would represent the number of years past 2000 (so that in the year 2040, t=40). pjp: does this help at all? Let us know where y ou stand and what you need help with to be able to complete this problem solving.

OpenStudy (whpalmer4):

After another 20 years, the population is multiplied by 0.98 again to reduce it by 2%. 3332*0.98 = (don't forget to round to the nearest whole number!) @mathmale It's unclear that the OP is really expected to understand an exponential growth/decay model. The problem asks for the population at a date where you can compute the answer just by multiplying by a percentage twice; I would expect any worthwhile exponential growth/decay model problem to force you to compute an exponential!